OpenStudy (jacksonjrb):

Find the directrix, focus, and vertex of the parabola whose equation is: y^2-8y+116=-20x

OpenStudy (freckles):

this involves completing the square you need to write this in the form \[(y-k)^2=4p(x-h)\]

OpenStudy (freckles):

so first let's look at the y^2-8y+? we need to figure out what ? should be so we can complete the square

OpenStudy (jacksonjrb):

16

OpenStudy (freckles):

perfect y^2-8y+16=(y-4)^2

OpenStudy (freckles):

\[y^2-8y+116=-20x \\ y^2-8y+16+100=-20x \\ (y-4)^2+100=-20x\] we also need to isolate the squared part

OpenStudy (freckles):

just in it is here: \[(y-k)^2=4p(x-h)\]

OpenStudy (freckles):

if you need a hint to do this.. let me know

OpenStudy (jacksonjrb):

\[(y-4)^2=-20(x-5)\]

OpenStudy (jacksonjrb):

x+5

OpenStudy (freckles):

cool

OpenStudy (freckles):

\[(y-4)^2=-20(x+5)\]

OpenStudy (jacksonjrb):

Vertex: (-5,4)

OpenStudy (freckles):

yep

OpenStudy (freckles):

now in place of 4p we see -20 so p is ?

OpenStudy (jacksonjrb):

-5

OpenStudy (freckles):

right |p| is the distance between the vertex and the directrix or the distance between the vertex and the focus (since the square is on the y stuff) also the - part means the parabola is opened to the left if it was + it would be opened to the right ------ (if the square was on the x stuff) - means down + means up

OpenStudy (jacksonjrb):

so the focus is (-5,9)?

OpenStudy (freckles):

you chanced the wrong coordinate